Hi i'm not sure this is the exact correct category for my specific word problem but this is it. Please show me your work if possible, I have no idea how to do this, and this would help greatly, Thank you:
12.Find four consecutive ODD integers such that the third is the sum of the fourth and twice the second?

Answer by Alan3354(69443)

12.Find four consecutive ODD integers such that the third is the sum of the fourth and twice the second?
-----------------
Integers are n, n+2, n+4 and n+6
n+4 = n+6 + 2(n+2)
n+4 = 3n + 10
-2n = 6
n = -3
------
-3, -1, +1, +3